OpenStudy (anonymous):

Expand log(subscript 2) 7x^3/y

OpenStudy (jdoe0001):

\(\bf log_2\left(\cfrac{7x^3}{y} \right)\) ?

OpenStudy (jdoe0001):

\(\bf \large \frac{log_2({7x^3})}{y}\) ?

OpenStudy (anonymous):

I'm sorry your first equation was correct.

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif look ^ at the rule 2

OpenStudy (anonymous):

I have \[\log_{2}7x^3\log_{2}y \]

OpenStudy (anonymous):

I am subtracting. Is there a way to simplify?

OpenStudy (jdoe0001):

look at rule 1 :)

OpenStudy (jdoe0001):

$$\bf log_2\left(\color{blue}{\frac{7x^3}{y}} \right)\\ \implies log_2(\color{blue}{7x^3})-log_2(y)\\ log_2(7)+log_2(\color{blue}{x^3})-log_2(y)\\ log_2(7)+3log_2(x)-log_2(y)\\ $$

OpenStudy (jdoe0001):

after rul 1, then rule 3 in the picture :)

OpenStudy (anonymous):

haha I'm beginning to see the answer :)

OpenStudy (jdoe0001):

well, that's the answer, it's just expanded

OpenStudy (anonymous):

. \[\log_{2}7+3\log_{2}(x)-\log_{2}(y) \]

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

I dont think that can be simplified any further :D I have 1 more problem but i dont see how it follows the rules of exponents.

OpenStudy (jdoe0001):

ok